OpenStudy (anonymous):

2(3x+6)=3(4x-2) solve pls

OpenStudy (bahrom7893):

6x + 12 = 12x - 6, divide everything by 6 x + 2=2x - 1 x-2x = -1-2 -x = -3 x=3

OpenStudy (bahrom7893):

Please click on good answer if I helped, thanks =)

OpenStudy (anonymous):

where will i click it? do u know how to check?

OpenStudy (bahrom7893):

plug in x =3 back into original: 2(3*3+6) = 3(4*3-2) 2(9+6) = 3(12-2) 2*15 = 3 * 10 30 = 30 so the answer is correct and you can click Good Answer on the upper right corner of my answer.. all the way to the right of group star.

OpenStudy (anonymous):

thank you so much for the help. but i have 1more. can you solve it again? 3(y-4)+3(y-2)

OpenStudy (bahrom7893):

is equal to what?

OpenStudy (anonymous):

3(y-4)+3(y-2)0

OpenStudy (anonymous):

3(y-4)=3(y-2)

OpenStudy (bahrom7893):

did you mean: 3(y-4)+3(y-2)=0, or 3(y-4)=3(y-2)

OpenStudy (anonymous):

3(y-4)=3(y-2) this one

OpenStudy (bahrom7893):

okay so: 3y - 12 = 3y - 6 3y-3y = -6 + 12 0 = 6, this is not true, so there are no solutions...

OpenStudy (anonymous):

ok. thank you